Assume that the number of pieces of junk mail per day that a person receives in their mail box follows the poisson distribution and averages 4.3 pieces per day. what is the probability that this person will receive more than two pieces of junk mail tomorrow?
Answer #1

Solution

using Poisson distribution

Average = λ = 4.3

P(X=x)=(e−λ⋅λx)/(x!)

P(X>2)=1−P(X≤2) = 1-(P(0)+P(1)+P(2))

= 1-(0.013568559012201+0.058344803752464 +0.1254413280678) = 1-0.19735469083246 = 0.80264530916754

Answer : 0.8026